# Machine learning a fixed point action for SU(3) gauge theory with a gauge equivariant convolutional neural network

### v1

@@ Abstract
Fixed point lattice actions are designed to have continuum classical properties unaffected by discretization effects and reduced lattice artifacts at the quantum level. They provide a possible way to extract continuum physics with coarser lattices, thereby allowing to circumvent problems with critical slowing down and topological freezing toward the continuum limit. A crucial ingredient for practical applications is to find an accurate and compact parametrization of a fixed point action, since many of its properties are only implicitly defined. Here we use machine learning methods to revisit the question of how to parametrize fixed point actions. In particular, we obtain a fixed point action for four-dimensional SU(3) gauge theory using convolutional neural networks with exact gauge invariance. The large operator space allows us to find superior parametrizations compared to previous studies, a necessary first step for future Monte Carlo simulations.
@@ I Introduction
Lattice regularization is the tool of choice to study nonperturbative properties of quantum field theories starting from first principles [1]. Modern lattice QCD simulations have attained a high level of precision and for some important Standard Model quantities, e.g., the QCD coupling at the electroweak scale $\alpha_{S}(\mu=m_{Z})$, they provide the current most accurate determination [2]. Increased precision has amplified systematic issues relevant to any lattice calculation, such as the extrapolation to the continuum limit. Numerical simulations become rapidly more costly as the lattice spacing is reduced to zero, not only due to the increased resolution at fixed physical volume, but also due to the increased autocorrelation times (critical slowing down) in generating statistically independent samples in Monte Carlo Markov chains and the related problem of suppressed tunneling between sectors of different topological charge (topological freezing[3]. For a robust continuum prediction, a range of lattice spacings is necessary, requiring a delicate balance between the control of discretization artifacts on coarse lattices on the one hand and the increased cost of simulating on finer lattices on the other.
@@
Here we propose to follow a complementary approach in order to solve both critical slowing down and topological freezing by using a lattice action with no or highly suppressed lattice artifacts. Such an action in principle allows simulations on very coarse lattices where both problems are absent, while at the same time lattice artifacts can be kept so small that a solid continuum limit can be taken. The advantage of this approach is the immediate applicability to gauge field theories in four dimensions without encountering scalability issues, once a highly improved action is found.
There is a long history of designing improved lattice actions to reduce discretization effects, bringing simulations at coarser lattice spacing into the scaling regime. One such program, Symanzik improvement [14, 15, 16, 17], removes lattice artifacts in some physical quantities order by order in the lattice spacing $a$. In a lattice gauge theory, this can be achieved, for example, by building a lattice action combining plaquettes and closed six-link loops. By construction, such an approach involves a perturbative expansion at weak coupling. A radically different approach makes use of renormalization group (RG) properties to design lattice actions where artifacts are removed completely to all orders. The construction of such quantum perfect actions is an extremely ambitious goal and is in general difficult to achieve. In asymptotically free theories, such as QCD, a constructive method can be designed based on the fixed point (FP) of RG transformations which yields lattice actions without lattice artifacts at the classical level, i.e., for on-shell quantities [18]. These so-called classically perfect actions, or FP actions in short, are in general expected to show suppressed lattice artifacts even at the quantum level. The FP action approach was used to study the O(3) nonlinear $\sigma$–model, SU($3$) pure gauge theory, and full QCD, with promising indications of much-reduced cutoff dependence in Monte Carlo simulations [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. However, the increased numerical cost of simulating FP gauge actions made it difficult at that time to draw firm conclusions on the level of improvement. Given the intervening dramatic increase in computing capability, this is no longer an obstacle and pushing the FP approach to higher accuracy has in principle become feasible.
@@
For any improved lattice action, the true test of how much lattice artifacts are reduced in the full quantum theory is only possible through Monte Carlo simulations. In this paper we focus on describing in detail the already challenging first step, to parametrize an FP action, in particular for the four-dimensional SU(3) gauge theory, using L-CNNs and ML techniques. This allows us to compare with previous studies of the FP parametrization and also serves as a proof of concept that ML can be accurately used in this task. The end result is that the very expressive nature of L-CNNs enables us to find a much more accurate parametrization of the SU($3$) FP action than previously possible. This success constitutes the first necessary step toward future Monte Carlo studies and ultimately toward the construction of a quantum perfect action without any lattice artifacts.
@@ II Theoretical setup
Figure 1: A sketch of the renormalization group flow and the renormalized trajectory (RT) in the infinite-dimensional coupling space, with the gauge coupling as the only relevant direction. The fixed point is on the critical surface $\beta\rightarrow\infty$ where $\xi/a=\infty$ for any physical scale $\xi$, with the values of the critical couplings $c^{\rm FP}_{n}$ determined by the specific form of the RG blocking. The FP action uses the same coupling values at finite $\beta$, tracking the RT very closely at weak coupling.
@@
For asymptotically free gauge theories, the only relevant coupling is the gauge coupling $g$ and the continuum is approached in the weak coupling limit $\beta\rightarrow\infty$. On the critical surface, where $\xi/a=\infty$, the irrelevant couplings ${c_{0},c_{1},\ldots}$ flow into a fixed point as shown in Figure 1. The FP couplings $\{c_{0}^{\text{FP}},c_{1}^{\text{FP}},\ldots\}$ are determined once the form of the RG blocking is prescribed. Slightly off the critical surface, the couplings first flow toward then away from the fixed point, approaching the renormalized trajectory (RT) which describes the flow starting from the FP in the relevant direction of the gauge coupling. Along the RT, the lattice theory is quantum perfect, with no lattice artifacts at all, because it is connected back to the continuum theory on the critical surface. The FP couplings define the so-called FP action ${\mathcal{A}}^{\rm FP}$. When it is used at finite values of $\beta$, it tracks the RT very closely at sufficiently weak coupling, cf. Figure 1. The FP action can be shown to be classically perfect [18, 20], i.e., it has no lattice artifacts of $O(a^{2n})$ to all orders on field configurations fulfilling the equations of motions. Artifacts of $O(g^{2}a^{2n})$ are, however, present but suppressed for small $g$.
@@
Figure 2: The leading couplings $\rho_{\mu\nu}(r)$ of the perturbative FP action, from [24]. The blocking kernel $T[U,V]$ is designed to maximize the exponential decay of the couplings, with $\exp(-3.4r)$ shown as a visual guide.
A crucial question concerns the locality of the FP action or, more generally, the action ${\mathcal{A}}^{\prime}$ in Eq. (1). In order to guarantee universality, the couplings must fall off exponentially in the separation between fields. One can design RGTs which force an exponential decay and find the one maximizing it, so that beyond some separation the couplings are small enough to be negligible and in practical applications can be omitted. Some guidance for a good choice of blocking kernel can be provided perturbatively [20, 19]. At weak coupling (and with some gauge fixing), the FP action can be expanded in terms of the gauge potential $A^{a}_{\mu}(x)$, only keeping terms up to quadratic order in the potential. The resulting action can be expressed in terms of couplings $\rho_{\mu\nu}(r)$ for fields $A_{\mu}(x)$ and $A_{\nu}(y)$ at separation $r=|x-y|$. An optimal choice of the blocking and the RGT w.r.t. locality was found in [24]. Figure 2 shows the corresponding largest perturbative couplings which fall off exponentially in magnitude with $\sim\exp(-3.4r)$. It is this RGT which we employ in our work and we give its details in Appendix D.
The FP action $\mathcal{A}^{\rm FP}$ and its properties are defined only implicitly through the FP Eq. (3), where $\mathcal{A}^{\rm FP}$ appears on both sides. The FP equation is therefore iterative: On the right-hand side of Eq. (3), the value of $\mathcal{A}^{\rm FP}[U]$ can be determined through a second minimization over even finer gauge configurations $U^{\prime}$, and so on, until we reach a configuration so smooth that any (reasonable) lattice discretization of the contimuum Yang-Mills gauge action can be used to calculate the inception value of the action. In practice, instead of iterating the FP equation, one can shortcut the procedure and, for sufficiently smooth fine configurations, make use of existing approximate parametrizations of $\mathcal{A}^{\rm FP}[U]$. Previous parametrizations include linear combinations of plaquette, rectangular, and parallelogram loops with various powers of their traces [24], or combinations of thin-link and smeared-link plaquette traces $u_{x,\mu\nu}$ and $w_{x,\mu\nu}$ with various powers of the form
@@ III Fixed point data
As described in the previous section, for practical reasons one employs a parametrization of $\mathcal{A}^{\rm FP}[U]$ for the minimization procedure and the question arises how this approximation affects the true value $\mathcal{A}^{\rm FP}[V]$. Since the RG blocking increases the lattice spacing by a factor of two in each RGT step, the action density on the fine configuration $U$ is at least a factor of 16 smaller than on the coarse configuration $V$, and in practice is even smaller, because of the sizable positive contribution from the blocking kernel $T[U,V]$.
@@
Another potential error on the FP data $\mathcal{A}^{\rm FP}[V]$ may originate from inaccurate minimization of the right-hand side of Eq. (3). The minimization on each configuration starts from an initial random fine configuration $U$ and then sequentially updates each link $U_{n,\mu}$ with an adaptive rotation in color space. Each iteration corresponds to a pass through the entire volume. We show two typical examples of this minimizing procedure in Fig. 4 for two coarse configurations $V$ on $8^{4}$ volume generated with $\beta_{\rm wil}=6.0$ (top plot) and $\beta_{\rm wil}=5.4$ (bottom plot). As shown in the figure, on smoother configurations at $\beta_{\rm wil}=6.0$ the minimization converges quickly, while on rougher configurations at $\beta_{\rm wil}=5.4$ as expected it takes somewhat longer to reach a similar level of convergence. In any case, we see from the illustrations that even in those cases, the error on the value of the FP action $\mathcal{A}^{\rm FP}[V]$ is negligible.
@@
Taking as input a fine instanton configuration with some chosen value of instanton radius $\rho$, we produce a coarse configuration $V$ using the RG blocking. If the topological properties are intact on the coarse side, the FP action should be unchanged by minimization, reproducing an instanton solution on the fine side. We see tests of this in Fig. 5. Starting from a fine instanton solution on a $16^{4}$ volume, the coarse $8^{4}$ configuration $V$ is produced via RG blocking and then fed into the minimization procedure. The upper plot is for an instanton originally of radius $\rho/a=3.0$; under minimization, the action is essentially unchanged, with a very small contribution from the blocking kernel $T[U,V]$, meaning the blocked configuration is also an instanton solution. The inset shows the rapid convergence of $\mathcal{A}[U]+T[U,V]$ in the minimization. The lower plot is for an instanton originally of radius $\rho/a=1.5$; once RG-blocked, the instanton is lost, as $T[U,V]$ becomes much larger during minimization and the minimized total action $\mathcal{A}[U]+T[U,V]$ is below the continuum value $4\pi^{2}$, i.e., the topological features are lost because the instanton can non longer be resolved at the level of the coarse lattice spacing $a^{\prime}=2a$. Note that with the RGT-III blocking employed in this work, instantons fall through the lattice for radii $\rho/a^{\prime}\lesssim 0.85$. In order to embrace this specific classical property of the FP action, we generate a set of coarse configurations through blocking fine instanton configurations with $\rho/a$ ranging from 1.1 to 3.0. The corresponding FP action values and derivatives provided by the minimization form part of the FP training data set. Further details of instanton solutions on the lattice are given in Appendix E.
@@ IV Machine learning model
Machine learning is being applied across a vast array of fields [34, 35, 36, 37, 38, 39]. Focused more specifically on lattice field theory, it has been used to in a range of topics, including the identification of phase transitions and their underlying critical exponents [40], the generation of decorrelated Markov chains through normalizing [9] or trivializing [12] flow transformations, inverting renormalization group transformations in spin models [41], the finite-temperature deconfinement phase transition in SU(2) and SU(3) pure gauge theory [42, 43], preconditioning of lattice Dirac operators [44, 45], and the connection between machine learning diffusion models and stochastic quantization of field theories through Langevin dynamics [46]. A recent review of some of this work is given in [47]. In our context, we need a tool to parametrize a lattice action in a highly general form, maintaining exact gauge invariance. The necessary architecture has already been developed in [30] with the lattice gauge equivariant convolutional neural network (L-CNN).
@@ IV.2 Parametrizing actions using L-CNNs
L-CNNs built from multiple bilinear convolutions with a final trace layer at the end of the network can be used to express a large class of gauge invariant scalar functions on the lattice. However, there are a few additional requirements to parametrize gauge invariant actions. The first requirement is a normalization condition: the output of a parametrization $\mathcal{A}^{\mathrm{L-CNN}}[V]$ must approach the Yang-Mills action $S_{\mathrm{YM}}[A_{\mu}(x)]/\beta$ for sufficiently smooth gauge configurations $V_{x,\mu}\approx\exp(iaA_{\mu}(x))$. Secondly, one may require that the naive continuum limit $a\rightarrow 0$ is reached in a particular way such that lattice artifacts of certain observables are suppressed to some desired order, along the lines of Symanzik improvement. A third condition is that the parametrized action should be positive for all gauge configurations. Finally, we require the action to be local, which means that the parametrization should be expressible as a sum over lattice sites of finite-length Wilson loops and their products.

### v2

@@ Abstract
Fixed point lattice actions are designed to have continuum classical properties unaffected by discretization effects and reduced lattice artifacts at the quantum level. They provide a possible way to extract continuum physics with coarser lattices, thereby allowing one to circumvent problems with critical slowing down and topological freezing toward the continuum limit. A crucial ingredient for practical applications is to find an accurate and compact parametrization of a fixed point action, since many of its properties are only implicitly defined. Here we use machine learning methods to revisit the question of how to parametrize fixed point actions. In particular, we obtain a fixed point action for four-dimensional SU(3) gauge theory using convolutional neural networks with exact gauge invariance. The large operator space allows us to find superior parametrizations compared to previous studies, a necessary first step for future Monte Carlo simulations and scaling studies.
@@ I Introduction
Lattice regularization is the tool of choice to study nonperturbative properties of quantum field theories starting from first principles [1]. Modern lattice QCD simulations have attained a high level of precision and for some important Standard Model quantities, e.g., the QCD coupling at the electroweak scale $\alpha_{S}(\mu=m_{Z})$, they provide the current most accurate determination [2]. Increased precision has amplified systematic issues relevant to any lattice calculation, such as the extrapolation to the continuum limit. Numerical simulations become rapidly more costly as the lattice spacing is reduced, not only due to the increased resolution at fixed physical volume, but also due to the increased autocorrelation times (critical slowing down) in generating statistically independent samples in Monte Carlo Markov chains and the related problem of suppressed tunneling between sectors of different topological charge (topological freezing[3]. For a robust continuum prediction, a range of lattice spacings is necessary, requiring a delicate balance between the control of discretization artifacts on coarse lattices on the one hand and the increased cost of simulating on finer lattices on the other.
@@
Here we propose to follow a complementary approach in order to solve both critical slowing down and topological freezing by using a lattice action with highly suppressed lattice artifacts. Such an action in principle allows simulations on very coarse lattices where both problems are absent, while at the same time lattice artifacts can be kept so small that a solid continuum limit can be taken. The advantage of this approach is the immediate applicability to gauge field theories in four dimensions without encountering scalability issues, once a highly improved action is found.
There is a long history of designing improved lattice actions to reduce discretization effects, bringing simulations at coarser lattice spacing into the scaling regime. One such program, Symanzik improvement [14, 15, 16, 17], removes lattice artifacts in some physical quantities order by order in the lattice spacing $a$. In a lattice gauge theory, this can be achieved, for example, by building a lattice action combining plaquettes and closed six-link loops. By construction, such an approach involves a perturbative expansion at weak coupling. A radically different approach makes use of renormalization group (RG) properties to design lattice actions where artifacts are removed completely to all orders. The construction of such quantum perfect actions is an extremely ambitious goal and is in general impossible to achieve. In asymptotically free theories, such as QCD, a constructive method can be designed based on the fixed point (FP) of RG transformations which yields lattice actions without lattice artifacts at the classical level, i.e., for on-shell quantities [18]. These so-called classically perfect actions, or FP actions in short, are in general expected to show suppressed lattice artifacts at sufficiently small gauge coupling $g$ even at the quantum level. The FP action approach was used to study the O(3) nonlinear $\sigma$–model, SU($3$) pure gauge theory, and full QCD, with promising indications of much-reduced cutoff dependence in Monte Carlo simulations [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. However, the increased numerical cost of simulating FP gauge actions made it difficult at that time to draw firm conclusions on the level of improvement. Given the intervening dramatic increase in computing capability, this is no longer an obstacle and pushing the FP approach to higher accuracy has in principle become feasible.
@@
For any improved lattice action, the true test of how much lattice artifacts are reduced in the full quantum theory is only possible through Monte Carlo simulations. In this paper we focus on describing in detail the already challenging first step, to parametrize an FP action, in particular for the four-dimensional SU(3) gauge theory, using L-CNNs and ML techniques. This allows us to compare with previous studies of the FP parametrization and also serves as a proof of concept that ML can be accurately used in this task. The end result is that the very expressive nature of L-CNNs enables us to find a much more accurate parametrization of the SU($3$) FP gauge action than previously possible. This conceptual success constitutes the first necessary step toward future Monte Carlo studies and ultimately toward the construction of a (approximate) quantum perfect action with strongly suppressed lattice artifacts.
@@ II Theoretical setup
Figure 1: A sketch of the renormalization group flow and the renormalized trajectory (RT) in the infinite-dimensional coupling space, with the gauge coupling as the only relevant direction. The fixed point is on the critical surface $\beta\rightarrow\infty$ where $\xi/a=\infty$ for any physical scale $\xi$, with the values of the critical couplings $c^{\rm FP}_{n}$ determined by the specific form of the RG blocking. The FP action uses the same coupling values at finite $\beta$, tracking the RT closely at sufficiently weak coupling.
@@
For asymptotically free gauge theories, the only relevant coupling is the gauge coupling $g$ and the continuum is approached in the weak coupling limit $\beta\rightarrow\infty$. On the critical surface, where $\xi/a=\infty$, the irrelevant couplings ${c_{0},c_{1},\ldots}$ flow into a fixed point as shown in Figure 1. The FP couplings $\{c_{0}^{\text{FP}},c_{1}^{\text{FP}},\ldots\}$ are determined once the form of the RG blocking is prescribed. Slightly off the critical surface, the couplings first flow toward then away from the fixed point, approaching the renormalized trajectory (RT) which describes the flow starting from the FP in the relevant direction of the gauge coupling. Along the RT, the lattice theory is quantum perfect, with no lattice artifacts at all, because it is connected back to the continuum theory on the critical surface. The FP couplings define the so-called FP action ${\mathcal{A}}^{\rm FP}$. When it is used at finite values of $\beta$, it tracks the RT closely at sufficiently weak coupling, cf. Figure 1. The FP action can be shown to be classically perfect [18, 20], i.e., it has no lattice artifacts of $O(a^{2n})$ to all orders on field configurations fulfilling the equations of motions. Artifacts of $O(g^{2}a\log a,g^{2}a^{2})$ are, however, present but suppressed for small $g$.
@@
Figure 2: The leading couplings $\rho_{\mu\nu}(r)$ of the perturbative FP action, from [24]. The blocking kernel $T[U,V]$ is designed to maximize the exponential decay of the couplings, with $\exp(-3.4r/a)$ shown as a visual guide.
A crucial question concerns the locality of the FP action or, more generally, the action ${\mathcal{A}}^{\prime}$ in Eq. (1). In order to guarantee universality, the couplings must fall off exponentially in the separation between fields. One can design RGTs which force an exponential decay and find the one maximizing it, so that beyond some separation the couplings are small enough to be negligible and in practical applications can be omitted. Some guidance for a good choice of blocking kernel can be provided perturbatively [20, 19]. At weak coupling (and with some gauge fixing), the FP action can be expanded in terms of the gauge potential $A^{a}_{\mu}(x)$, only keeping terms up to quadratic order in the potential. The resulting action can be expressed in terms of couplings $\rho_{\mu\nu}(r)$ for fields $A_{\mu}(x)$ and $A_{\nu}(y)$ at separation $r=|x-y|$. An optimal choice of the blocking and the RGT w.r.t. locality was found in [24]. Figure 2 shows the corresponding largest perturbative couplings which fall off exponentially in magnitude with $\sim\exp(-3.4r/a)$. It is this RGT which we employ in our work and we give its details in Appendix D.
The FP action $\mathcal{A}^{\rm FP}$ and its properties are defined only implicitly through the FP Eq. (3), where $\mathcal{A}^{\rm FP}$ appears on both sides. The FP equation is therefore iterative: On the right-hand side of Eq. (3), the value of $\mathcal{A}^{\rm FP}[U]$ can be determined through a second minimization over even finer gauge configurations $U^{\prime}$, and so on, until we reach a configuration so smooth that any (reasonable) lattice discretization of the continuum Yang-Mills gauge action can be used to calculate the inception value of the action. In practice, instead of iterating the FP equation, one can shortcut the procedure and, for sufficiently smooth fine configurations, make use of existing approximate parametrizations of $\mathcal{A}^{\rm FP}[U]$. Previous parametrizations include linear combinations of plaquette, rectangular, and parallelogram loops with various powers of their traces [24], or combinations of thin-link and smeared-link plaquette traces $u_{x,\mu\nu}$ and $w_{x,\mu\nu}$ with various powers of the form
@@ III Fixed point data
As described in the previous section, for practical reasons one employs a parametrization of $\mathcal{A}^{\rm FP}[U]$ for the minimization procedure and the question arises how this approximation affects the true value $\mathcal{A}^{\rm FP}[V]$. Since the RG blocking increases the lattice spacing by a factor of 2 in each RGT step, the action density on the fine configuration $U$ is at least a factor of 16 smaller than on the coarse configuration $V$, and in practice is even smaller, because of the sizable positive contribution from the blocking kernel $T[U,V]$.
@@
Another potential error on the FP data $\mathcal{A}^{\rm FP}[V]$ may originate from inaccurate minimization of the right-hand side of Eq. (3). The minimization on each configuration starts from an initial random fine configuration $U$ and then sequentially updates each link $U_{n,\mu}$ with an adaptive rotation in color space. Each iteration corresponds to a pass through the entire volume. We show two typical examples of this minimizing procedure in Fig. 4 for two coarse configurations $V$ on $8^{4}$ volume generated with $\beta_{\rm wil}=6.0$ (top plot) and $\beta_{\rm wil}=5.4$ (bottom plot). As shown in the figure, on smoother configurations at $\beta_{\rm wil}=6.0$ the minimization converges quickly, while on rougher configurations at $\beta_{\rm wil}=5.4$, as expected, it takes somewhat longer to reach a similar level of convergence. In any case, we see from the illustrations that even in those cases, the error on the value of the FP action $\mathcal{A}^{\rm FP}[V]$ is negligible.
@@
Taking as input a fine instanton configuration with some chosen value of instanton radius $\rho$, we produce a coarse configuration $V$ using the RG blocking. If the topological properties are intact on the coarse side, the FP action should be unchanged by minimization, reproducing an instanton solution on the fine side. We see tests of this in Fig. 5. Starting from a fine instanton solution on a $16^{4}$ volume, the coarse $8^{4}$ configuration $V$ is produced via RG blocking and then fed into the minimization procedure. The upper plot is for an instanton originally of radius $\rho/a=3.0$; under minimization, the action is essentially unchanged, with a very small contribution from the blocking kernel $T[U,V]$, meaning the blocked configuration is also an instanton solution. The inset shows the rapid convergence of $\mathcal{A}[U]+T[U,V]$ in the minimization. The lower plot is for an instanton originally of radius $\rho/a=1.5$; once RG-blocked, the instanton is lost, as $T[U,V]$ becomes much larger during minimization and the minimized total action $\mathcal{A}[U]+T[U,V]$ is below the continuum value $4\pi^{2}$, i.e., the topological features are lost because the instanton can no longer be resolved at the level of the coarse lattice spacing $a^{\prime}=2a$. Note that with the RGT-III blocking employed in this work, instantons fall through the lattice for radii $\rho/a^{\prime}\lesssim 0.85$. In order to embrace this specific classical property of the FP action, we generate a set of coarse configurations through blocking fine instanton configurations with $\rho/a$ ranging from 1.1 to 3.0. The corresponding FP action values and derivatives provided by the minimization form part of the FP training data set. Further details of instanton solutions on the lattice are given in Appendix E.
@@ IV Machine learning model
Machine learning is being applied across a vast array of fields [34, 35, 36, 37, 38, 39]. Focused more specifically on lattice field theory, it has been used in a range of topics, including the identification of phase transitions and their underlying critical exponents [40], the generation of decorrelated Markov chains through normalizing [9] or trivializing [12] flow transformations, inverting renormalization group transformations in scalar field theory [41], the finite-temperature deconfinement phase transition in SU(2) and SU(3) pure gauge theory [42, 43], preconditioning of lattice Dirac operators [44, 45], and the connection between machine learning diffusion models and stochastic quantization of field theories through Langevin dynamics [46]. A recent review of some of this work is given in [47]. In our context, we need a tool to parametrize a lattice action in a highly general form, maintaining exact gauge invariance. The necessary architecture has already been developed in [30] with the lattice gauge equivariant convolutional neural network (L-CNN).
@@ IV.2 Parametrizing actions using L-CNNs
L-CNNs built from multiple bilinear convolutions with a final trace layer at the end of the network can be used to express a large class of gauge invariant scalar functions on the lattice. However, there are a few additional requirements to parametrize gauge invariant actions. The first requirement is a normalization condition: the output of a parametrization $\mathcal{A}^{\textrm{L-CNN}}[V]$ must approach the Yang-Mills action $S_{\mathrm{YM}}[A_{\mu}(x)]/\beta$ for sufficiently smooth gauge configurations $V_{x,\mu}\approx\exp(iaA_{\mu}(x))$. Secondly, one may require that the naive continuum limit $a\rightarrow 0$ is reached in a particular way such that lattice artifacts of certain observables are suppressed to some desired order, along the lines of Symanzik improvement. A third condition is that the parametrized action should be positive for all gauge configurations. Finally, we require the action to be local, which means that the parametrization should be expressible as a sum over lattice sites of finite-length Wilson loops and their products.